# Dynamic asymptotic dimension growth for group actions and groupoids

### v1

@@ 4.2. Example: coarse spaces with asymptotic dimension growth
In this subsection, we will discuss the connection between our notion and coarse geometry. More specifically, for a discrete metric space of bounded geometry $X$, we will show that the asymptotic dimension growth for $X$ is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid $G(X)$.
@@
    equip $G(X)$ with the weak topology coming from this union, this is a locally compact topology for which each $\overline{E_{R}}$ is compact open subset of $G(X)$;
@@
The coarse groupoid $G(X)$ is $\sigma$-compact, locally compact Hausdorff, principal and étale. Recall from [MW20] that there is a canonical length function $\ell$ on $G(X)$ defined by extending the metric $d:X\times X\rightarrow[0,\infty]$ to $\beta(X\times X)$ and observing that it takes finite values on $G(X)$. The canonical length function $\ell$ on $G(X)$ has the following properties:
@@
    $\ell$ is coarse. Since $X\times X$ is dense in $\beta(X\times X)$, for any $R>0$, the open set $\ell^{-1}([0,R))$ is contained in $\overline{\ell^{-1}([0,R))\cap X\times X}$, which is contained in the compact set $\overline{E_{R}}$. Hence, $\ell^{-1}([0,R))$ is precompact, and $\ell$ is proper. Observe that for any precompact subset $K$ of $G(X)$, $K$ must contained in a controlled set $E_{R}$ for some $R$ by the definition of topology of $G(X)$. Thus $\ell$ is controlled.
@@
In the case of coarse groupoids, the following theorem bridges the asymptotic dimension growth with the dynamic asymptotic dimension growth.
@@
Assume that $G(X)$ has dynamic asymptotic dimension function $\textnormal{dad}_{G(X)}$, and $X$ has dynamic asymptotic dimension function $\widecheck{\textnormal{ad}}_{X}$ as defined in Lemma 2.15. Below, we show that $\widecheck{\textnormal{ad}}_{X}\approx\textnormal{dad}_{G(X)}$. First, we need to verify that $\widecheck{\textnormal{ad}}_{X}\preceq\textnormal{dad}_{G(X)}$. For any $R>0$, for the open precompact subset $K=\{z\in G(X)\mid\ell(z)<R\}$ of $G(X)$, then there exists a controlled set $E_{R-\frac{1}{n}}$ for some positive integer $n$ such that $\overline{E_{R-\frac{1}{n}}}\subseteq K$. Observe that $\overline{E_{R-\frac{1}{n}}}$ is a compact open subset of $G(X)$ containing the unit space $G(X)^{(0)}=\beta X$. Let $U_{0},\cdots,U_{\textnormal{dad}_{G(X)}(R)}$ be an open cover of $G(X)^{(0)}=\beta X$ such that for each $i\in\{0,\cdots,\textnormal{dad}_{G(X)}(R)\}$, the set
@@
such that each $\mathcal{U}_{i}$ is $E_{R}$-separated. For each $i$, set
@@
is contained in $\overline{\mathcal{G}_{i}}$, a compact subgroupoid of $G(X)$ contained in the compact subset $\overline{F}$. Therefore, $\textnormal{dad}_{G(X)}(R)\leq\widecheck{\textnormal{ad}}_{X}(R)$, which yields that $\textnormal{dad}_{G(X)}\preceq\widecheck{\textnormal{ad}}_{X}$. Hence, we obtain $\textnormal{dad}_{G(X)}\approx\widecheck{\textnormal{ad}}_{X}$ as desired.
@@
Combining Theorem 4.12 and Theorem 4.14, we have the following result:
@@
The following example follows from Theorem 4.12 and Corollary 4.15.
@@
Let $X$ be a geodesic uniformly locally finite coarse median space with finite rank and at most exponential volume growth, then its corresponding coarse groupoid $G(X)$ has subexponential dynamic asymptotic dimension growth, and it is amenable.
@@ 4.3. Equivalence with asymptotic dimension growth
In this subsection, we will build the equivalence between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for group actions by resorting to the BLR condition. Using this result, we can obtain that countable discrete groups with subexponential asymptotic dimension growth all admit free, amenable, minimal actions on the Cantor set which have at most subexponential dynamic asymptotic dimension growth.
@@
is contained in the set in line 4.2, and is thus finite. This shows that $\textnormal{dad}_{\Gamma\curvearrowright\beta\Gamma}(R)\leq\textnormal{dad}_{\Gamma\curvearrowright X}(R)$, and thus $\textnormal{dad}_{\Gamma\curvearrowright\beta\Gamma}\leq\textnormal{dad}_{\Gamma\curvearrowright X}\preceq h$.
$(3)\Rightarrow(4)$: the transformation groupoid $\beta\Gamma\rtimes\Gamma$ has dynamic asymptotic dimension function $\textnormal{dad}_{\beta\Gamma\rtimes\Gamma}\preceq h$ by Lemma 4.9. On the other hand, $\beta\Gamma\rtimes\Gamma=G(\Gamma)$ by [Proposition 3.4, [STY02]], thus $G(\Gamma)$ has asymptotic dimension function $\textnormal{dad}_{G(X)}\preceq h$. By Theorem 4.12, $\Gamma$ has asymptotic dimension function $\textnormal{ad}_{X}\preceq h$.
$(4)\Rightarrow(5)$: Combining Theorem 4.14 and the assumption that $\Gamma$ has at most subexponential asymptotic dimension growth, we have that $\Gamma$ has property A. Then the canonical action of $\Gamma$ on $\beta\Gamma$ is amenable by [Theorem 3.3, [HR00]]. Further, given $R>0$, the finite subset $E=\{\gamma\in\Gamma\mid\ell(\gamma)<R\}$ and $\varepsilon>0$, there exists a continuous $f:\beta\Gamma\rightarrow\textnormal{Prob}_{\textnormal{ad}_{X}(R)}(\Gamma)$ such that
@@
Combining Theorem 4.17 with the results of Rørdam and Sierakowski in [Section 6, [RS12]], we obtain the following theorem.
@@
Step 1. In the first step, we will construct countable sets $\{V_{n}\}_{n\in\mathbb{N}}$. From the fact that $(4)$ implies $(5)$ of Theorem 4.17, we see that the action of $\Gamma$ on $\beta\Gamma$ is amenable and $\textnormal{ad}_{X}$-BLR for the family $\mathcal{F}$ of finite subgroups of $\Gamma$ and $\textnormal{ad}_{X}\approx h$. For each $n\in\mathbb{N}$, let $E_{n}=\{\gamma\in\Gamma\mid\ell(\gamma)<n\}$, which is a finite subset of $\Gamma$. By the properties of the length function $\ell$, there is a nested sequence
@@
such that $\Gamma=\bigcup\limits_{n}E_{n}$. Then for each $n\in\mathbb{N}$, there exist a $(\Gamma,\mathcal{F})$-complex $C_{n}$ of dimension at most $\textnormal{ad}_{X}(n)$ and a continuous $(E_{n},\frac{1}{n})$-equivariant map $f_{n}:\beta\Gamma\rightarrow C_{n}$. Since $\beta\Gamma$ is compact, we assume by Lemma 3.9 that $f_{n}(\beta\Gamma)\subset\textnormal{Prob}(S_{n})\cap C_{n}$, where $S_{n}$ is a finite subset of $C_{n}$ with cardinality $|S_{n}|\leq\textnormal{ad}_{X}(n)+1$. Further, we replace $C_{n}$ by $C_{n}\cap\textnormal{Prob}(\Gamma\cdot S_{n})$, we may assume that there are only finitely many $\Gamma$-orbits of $S_{n}$, which implies that the vertex set $V_{n}$ of $C_{n}$ is countable.
Step 2. In the next step, we show that $\Gamma$ admits a free and $\textnormal{ad}_{X}$-BLR action on a totally disconnected, metrizable compact space $\hat{A}$. Let $A$ be a unital $\Gamma$-invariant $C^{*}$-subalgebra of $\ell^{\infty}(\Gamma)$ , and let $\hat{A}$ denote the spectrum of $A$. By Gelfand duality, $\hat{A}$ identifies with a quotient space of $\beta\Gamma$ such that the quotient map $q:\beta\Gamma\rightarrow\hat{A}$ is equivariant.
@@
where each $t_{v}:\beta\Gamma\rightarrow[0,1]$ is a continuous map. Define $\Omega_{1}:=\{t_{v}\mid v\in V_{n},n\in\mathbb{N}\}$, which is a countable subset of $\ell^{\infty}(\Gamma)$ as $\ell^{\infty}(\Gamma)$ identifies with $C(\beta\Gamma)$. If $A$ contains $\Omega_{1}$, then the action of $\Gamma$ on $\hat{A}$ is $\textnormal{ad}_{X}$-BLR. Indeed, the maps $t_{v}:\beta\Gamma\rightarrow[0,1]$ descend to the maps $\hat{A}\rightarrow[0,1]$, whence the maps $f_{n}:\beta\Gamma\rightarrow C_{n}$ descend to the maps $f_{n}:\hat{A}\rightarrow C_{n}$ by [Lemma 3.5, [Hig00]]. On the other hand, [Lemma 6.4, [RS12]] shows that there exists a countable subset $\Omega_{2}$ of $\ell^{\infty}(\Gamma)$ such that if $A$ contains $\Omega_{2}$, then the action of $\Gamma$ on $\hat{A}$ is free.
Let $\Omega=\Omega_{1}\cup\Omega_{2}$, a countable subset of $\ell^{\infty}(\Gamma)$. According to [Lemma 6.7, [RS12]], there exists a countable $\Gamma$-invariant collection of projections $\mathcal{P}$ such that the $C^{*}$-algebra $C^{*}(\mathcal{P})$ generated by $\mathcal{P}$ contains $\Omega$. Let $A=C^{*}(\mathcal{P},1_{\ell^{\infty}(\Gamma)})$, then $\hat{A}$ is a $\Gamma$-space, and the action of $\Gamma$ is free and $\textnormal{ad}_{X}$-BLR. Since $A$ is unital and generated by countably many projections, the space $\hat{A}$ is compact, metrizible and totally disconnected.
@@
Compactness of $Y$ implies that any descending chain for this order has a non-empty intersection. Zorn’s lemma yields that there is a minimal element $X$. Thus the induced $\Gamma$ action on $X$ is minimal. Hence, $\Gamma$ admits a free, minimal, $\textnormal{ad}_{X}$-BLR action on a totally disconnected, metrizable compact space $X$. By Theorem 3.14, this action has dynamic asymptotic dimension function $\textnormal{dad}_{\Gamma\curvearrowright X}\leq\textnormal{ad}_{X}\preceq h$, and thus has dynamic asymptotic dimension growth at most $h$.
Step 4. In the final step, we finish the proof by showing that this $X$ is a copy of the Cantor set, and it thus suffices to show that it has no isolated points. Observe that if $x\in X$ is isolated, then the orbit of $\Gamma\cdot x$ is open, and is dense in $X$ by minimality, thus $X\backslash\Gamma\cdot x=\varnothing$. Since the action is free and $\Gamma$ is infinite, the open cover $\{\{\gamma x\}\mid\gamma\in\Gamma\}$ of $X=\Gamma\cdot x$ has no finite subcover which contradicts compactness. It follows that there is no isolated point. ∎

### v2

@@ 4.2. Example: coarse spaces with asymptotic dimension growth
In this subsection, we study a geometric class of groupoids. Given a discrete metric space of bounded geometry $X$, Skandalis, Tu, and Yu introduced an étale groupoid $G(X)$ in [STY02], called the coarse groupoid since it encodes many coarse geometric properties of the underlying space $X$. Moreover, we show that the asymptotic dimension growth for $X$ is equivalent to the dynamic asymptotic dimension growth for its associated coarse groupoid $G(X)$. This example served as our primary motivation and also inspired the terminology ‘dynamic asymptotic dimension growth’.
@@
    equip $G(X)$ with the weak topology coming from this union, this is a locally compact topology for which each $\overline{E_{R}}$ is a compact open subset of $G(X)$;
@@
The coarse groupoid $G(X)$ is $\sigma$-compact, locally compact Hausdorff, principal and étale [STY02]. Here a discrete metric space of bounded geometry is also called a uniformly locally finite metric space. Recall from [MW20] that there is a canonical length function $\ell$ on $G(X)$ defined by extending the metric $d:X\times X\rightarrow[0,\infty]$ to $\beta(X\times X)$ and observing that it takes finite values on $G(X)$. The canonical length function $\ell$ on $G(X)$ has the following properties:
@@
    $\ell$ is coarse. Since $X\times X$ is dense in $\beta(X\times X)$, for any $R>0$, the open set $\ell^{-1}([0,R))$ is contained in $\overline{\ell^{-1}([0,R))\cap X\times X}$, which is contained in the compact set $\overline{E_{R}}$. Hence, $\ell^{-1}([0,R))$ is precompact, which yields that $\ell$ is proper. Since $\ell$ is continuous, it is controlled.
@@
The following theorem bridges the dynamic asymptotic dimension growth for the coarse groupoid $G(X)$ with the asymptotic dimension growth for the underlying space $X$, which generalizes the finite case of [[GWY17], Theorem 6.4].
@@
Let us assume that $G(X)$ has dynamic asymptotic dimension function $\textnormal{dad}_{G(X)}$, and $X$ has asymptotic dimension function $\widecheck{\textnormal{ad}}_{X}$ as defined in Lemma 2.13. Below, we show that $\widecheck{\textnormal{ad}}_{X}\approx\textnormal{dad}_{G(X)}$. First, we need to verify that $\widecheck{\textnormal{ad}}_{X}\preceq\textnormal{dad}_{G(X)}$. For any $R>0$, for the open precompact subset $K=\{z\in G(X)\mid\ell(z)<R\}$ of $G(X)$, there exists a controlled set $E_{R-\frac{1}{n}}$ for some positive integer $n$ such that $\overline{E_{R-\frac{1}{n}}}\subseteq K$. Observe that $\overline{E_{R-\frac{1}{n}}}$ is a compact open subset of $G(X)$ containing the unit space $G(X)^{(0)}=\beta X$. Let $U_{0},\cdots,U_{\textnormal{dad}_{G(X)}(R)}$ be an open cover of $G(X)^{(0)}=\beta X$ such that for each $i\in\{0,\cdots,\textnormal{dad}_{G(X)}(R)\}$, the set
@@
and each $\mathcal{U}_{i}$ is $E_{R}$-separated. For each $i$, set
@@
is contained in $\overline{\mathcal{G}_{i}}$, a compact subgroupoid of $G(X)$ contained in the compact subset $\overline{F}$. Thus the subset $\{g\in K\mid s(g),r(g)\in U_{i}\}$ is contained in a compact open subgroupoid of $G(X)$. Therefore, $\textnormal{dad}_{G(X)}(R)\leq\widecheck{\textnormal{ad}}_{X}(R)$, which yields that $\textnormal{dad}_{G(X)}\preceq\widecheck{\textnormal{ad}}_{X}$. Hence, we obtain $\textnormal{dad}_{G(X)}\approx\widecheck{\textnormal{ad}}_{X}$ as desired.
@@
Since property A for $X$ is equivalent to the amenability of the corresponding coarse groupoid $G(X)$ [[STY02], Theorem 5.3]. Combining Theorems 4.16 and 4.18, we have the following result:
@@
The following example follows from Example 2.7, Theorem 4.16, and Corollary 4.19.
@@
Let $X$ be a geodesic uniformly locally finite coarse median space with finite rank and at most exponential volume growth. Then its corresponding coarse groupoid $G(X)$ has subexponential dynamic asymptotic dimension growth, and it is amenable.
@@ 4.3. Equivalence with asymptotic dimension growth
In this subsection, we will establish the equivalence between the asymptotic dimension growth for countable discrete groups and the dynamic asymptotic dimension growth for their group actions. As a result, we demonstrate that countable discrete groups with subexponential asymptotic dimension growth all admit free, amenable, minimal actions on the Cantor set which have at most subexponential dynamic asymptotic dimension growth.
Since property A for a countable discrete group $\Gamma$ is equivalent to the (topological) amenability of the natural action of $\Gamma$ on its Stone-Čech compactification $\beta\Gamma$ [HR00, Wil09]. A combination of Theorems 3.15, 4.16, 4.18 and Lemma 4.12 yields the following equivalences.
@@
is contained in the set in line (4.2), and is thus finite. This shows that $\textnormal{dad}_{\Gamma\curvearrowright\beta\Gamma}(R)\leq\textnormal{dad}_{\Gamma\curvearrowright X}(R)$, and thus $\textnormal{dad}_{\Gamma\curvearrowright\beta\Gamma}\leq\textnormal{dad}_{\Gamma\curvearrowright X}\preceq h$.
$(3)\Rightarrow(4)$: the transformation groupoid $\beta\Gamma\rtimes\Gamma$ has dynamic asymptotic dimension function $\textnormal{dad}_{\beta\Gamma\rtimes\Gamma}\preceq h$ by Lemma 4.12. On the other hand, $\beta\Gamma\rtimes\Gamma=G(\Gamma)$ by [[STY02], Proposition 3.4], thus $G(\Gamma)$ has dynamic asymptotic dimension function $\textnormal{dad}_{G(X)}\preceq h$. By Theorem 4.16, $\Gamma$ has asymptotic dimension function $\textnormal{ad}_{X}\preceq h$.
$(4)\Rightarrow(5)$: Combining Theorem 4.18 with the assumption that $\Gamma$ has at most subexponential asymptotic dimension growth, we have that $\Gamma$ has property A. Then the canonical action of $\Gamma$ on $\beta\Gamma$ is amenable by [[HR00], Theorem 3.3]. Further, given $R>0$, the finite subset $E=\{\gamma\in\Gamma\mid\ell(\gamma)<R\}$ and $\varepsilon>0$, there exists a continuous $f:\beta\Gamma\rightarrow\textnormal{Prob}_{\textnormal{ad}_{X}(R)}(\Gamma)$ such that
@@
Combining Theorem 4.21 with the results of Rørdam and Sierakowski in [[RS12], Section 6], we obtain the following theorem.
@@
Step 1. In the first step, we will construct countable sets $\{V_{n}\}_{n\in\mathbb{N}}$. From the fact that the statement $(4)$ implies the statement $(5)$ of Theorem 4.21, we see that the action of $\Gamma$ on $\beta\Gamma$ is amenable and is $\textnormal{ad}_{X}$-BLR for the family $\mathcal{F}$ of finite subgroups of $\Gamma$ with $\textnormal{ad}_{X}\approx h$. For each $n\in\mathbb{N}$, let $E_{n}=\{\gamma\in\Gamma\mid\ell(\gamma)<n\}$, which is a finite subset of $\Gamma$. By the properties of the length function $\ell$, there is a nested sequence
@@
such that $\Gamma=\bigcup\limits_{n}E_{n}$. Then for each $n\in\mathbb{N}$, there exist a $(\Gamma,\mathcal{F})$-complex $C_{n}$ of dimension at most $\textnormal{ad}_{X}(n)$ and a continuous $(E_{n},\frac{1}{n})$-equivariant map $f_{n}:\beta\Gamma\rightarrow C_{n}$. Since $\beta\Gamma$ is compact, we assume by Lemma 3.10 that $f_{n}(\beta\Gamma)\subseteq\textnormal{Prob}(S_{n})\cap C_{n}$, where $S_{n}$ is a finite subset of $C_{n}$ with cardinality $|S_{n}|\leq\textnormal{ad}_{X}(n)+1$. Further, we replace $C_{n}$ by $C_{n}\cap\textnormal{Prob}(\Gamma\cdot S_{n})$, we may assume that there are only finitely many $\Gamma$-orbits of $S_{n}$, which implies that the vertex set $V_{n}$ of $C_{n}$ is countable.
Step 2. In the next step, we show that $\Gamma$ admits a free, $\textnormal{ad}_{X}$-BLR action on a totally disconnected, metrizable compact space $\hat{A}$. Let $A$ be a unital $\Gamma$-invariant $C^{*}$-subalgebra of $\ell^{\infty}(\Gamma)$ , and let $\hat{A}$ denote the spectrum of $A$. By Gelfand duality, $\hat{A}$ identifies with a quotient space of $\beta\Gamma$ such that the quotient map $q:\beta\Gamma\rightarrow\hat{A}$ is equivariant.
@@
where each $t_{v}:\beta\Gamma\rightarrow[0,1]$ is a continuous map. Define $\Omega_{1}:=\{t_{v}\mid v\in V_{n},n\in\mathbb{N}\}$, which is a countable subset of $\ell^{\infty}(\Gamma)$ as $\ell^{\infty}(\Gamma)$ identifies with $C(\beta\Gamma)$. If $A$ contains $\Omega_{1}$, then the action of $\Gamma$ on $\hat{A}$ is $\textnormal{ad}_{X}$-BLR. Indeed, the maps $t_{v}:\beta\Gamma\rightarrow[0,1]$ descend to the maps $\hat{A}\rightarrow[0,1]$, whence the maps $f_{n}:\beta\Gamma\rightarrow C_{n}$ descend to the maps $f_{n}:\hat{A}\rightarrow C_{n}$ by [[Hig00], Lemma 3.5]. On the other hand, [[RS12], Lemma 6.4] shows that there exists a countable subset $\Omega_{2}$ of $\ell^{\infty}(\Gamma)$ such that if $A$ contains $\Omega_{2}$, then the action of $\Gamma$ on $\hat{A}$ is free.
Let $\Omega=\Omega_{1}\cup\Omega_{2}$, a countable subset of $\ell^{\infty}(\Gamma)$. According to [[RS12], Lemma 6.7], there exists a countable $\Gamma$-invariant collection of projections $\mathcal{P}$ such that the $C^{*}$-algebra $C^{*}(\mathcal{P})$ generated by $\mathcal{P}$ contains $\Omega$. Let $A=C^{*}(\mathcal{P},1_{\ell^{\infty}(\Gamma)})$, then $\hat{A}$ is a $\Gamma$-space, and the action of $\Gamma$ is free and $\textnormal{ad}_{X}$-BLR. Since $A$ is unital and generated by countably many projections, the space $\hat{A}$ is compact, metrizible and totally disconnected.
@@
Compactness of $Y$ implies that any descending chain for this order has a non-empty intersection. Zorn’s lemma yields that there is a minimal element $X$. Thus the induced $\Gamma$ action on $X$ is minimal. Hence, $\Gamma$ admits a free, minimal, $\textnormal{ad}_{X}$-BLR action on a totally disconnected, metrizable compact space $X$. By Theorem 3.15, this action has dynamic asymptotic dimension function $\textnormal{dad}_{\Gamma\curvearrowright X}\leq\textnormal{ad}_{X}\approx h$, and thus has dynamic asymptotic dimension growth at most $h$.
Step 4. In the final step, we finish the proof by showing that this $X$ is a copy of the Cantor set, and it thus suffices to show that it has no isolated points. Observe that if $x\in X$ is isolated, then the orbit of $\Gamma\cdot x$ is open, and is dense in $X$ by minimality, thus $X\backslash\Gamma\cdot x=\varnothing$. Since the action is free and $\Gamma$ is infinite, the open cover $\{\{\gamma x\}\mid\gamma\in\Gamma\}$ of $X=\Gamma\cdot x$ has no finite subcover, which contradicts compactness. It follows that there is no isolated point. ∎
Remark 4.23.
Suppose a countable discrete group $\Gamma$ has infinite asymptotic dimension and subexponential asymptotic dimension growth. In that case, it admits a free, minimal action on the Cantor set, and the dynamic asymptotic dimension of this action is infinite by [[Bon24], Theorem B].